OpenStudy (anonymous):

Solve for s: 5t-8w/s = 3w

OpenStudy (goldphenoix):

\[\large \large 5t - \frac{ 8w }{ s } = 3w\] Does it look like that?

OpenStudy (sumi29):

Is t supposed to be a parameter, or just a variable? If its simply a variable, then can't you just perform basic arithmetic to solve for s?

OpenStudy (anonymous):

Its like 5t-8w -------- = 30w s

OpenStudy (goldphenoix):

\[\large \large \frac{ 5t-8w }{s } = 3w\] Like that?

OpenStudy (anonymous):

yes im sort of new to this website idk how to work out everything just yet!

OpenStudy (goldphenoix):

Alright. So we want to find s. We want to get rid of the fraction. So we want to multiply s by 3w. What is 3w * s give you?

OpenStudy (anonymous):

3ws

OpenStudy (goldphenoix):

Good.  Remember we want to solve s, we must have s by itself.

OpenStudy (goldphenoix):

Now how can we get s by itself?

OpenStudy (anonymous):

Divide s into 8w and 3ws? to get it by itself

OpenStudy (goldphenoix):

Hmm. I'm not sure what you are trying to say. But we have 3ws on the right side. We just want s on the right side. So how can we get to 3ws to s?

OpenStudy (anonymous):

Subtract?

OpenStudy (goldphenoix):

Nope. What's the reciprocal of multiplication? Since 3w is a coefficient. (3 * w)

OpenStudy (anonymous):

s or 5?

OpenStudy (goldphenoix):

What do you mean?

OpenStudy (goldphenoix):

Do you know what's the reciprocal of multiplication?

OpenStudy (goldphenoix):

If you are not sure, then say it. :)

OpenStudy (uri):

Is that 8 or s?

OpenStudy (goldphenoix):

o.o???

OpenStudy (anonymous):

no idk

OpenStudy (uri):

Apsy he pooch rahi hun bhai

OpenStudy (uri):